Find the level of confidence assigned to an interval estimate of the mean formed using the following intervals. (Round your answers to four decimal places.)

(a) x − 0.99·σx to x + 0.99·σx


(b) x − 1.69·σx to x + 1.69·σx


(c) x − 2.22·σx to x + 2.22·σx


(d) x − 2.66·σx to x + 2.66·σx
Answer #1

Ans:

a)

confidence level,c=1-2*P(z<-0.99)

=1-2*normsdist(-0.99)=0.6778

b)

confidence level,c=1-2*P(z<-1.69)

=1-2*normsdist(-1.69)=0.9090

c)

confidence level,c=1-2*P(z<-2.22)

=1-2*normsdist(-2.22)=0.9736

d)

confidence level,c=1-2*P(z<-2.66)

=1-2*normsdist(-2.66)=0.9922